simplify:
1. 4x-2y-6x+4y
2. 6-3x-9-4x
3. 12x-2-6x-6
find x:
1. 4x-3=3x
2. 3x-2-x=4
3. 3x=6

Answer by ReadingBoosters(3246)

Simplify
1.)4x-2y-6x+4y
Combine like terms
4x-6x-2y+4y
-2x+2y
...
2.)6-3x-9-4x
Combine like terms
-3x-4x+6-9
-7x-3
...
3.)12x-2-6x-6
Combine like terms
12x-6x-2-6
6x-8
...
Solve for x
1). 4x-3=3x
Isolate x
4x-3x=3
x=3
...
2). 3x-2-x=4
Isolate x
3x-x=4+2
2x=6
Divide both sides by 2
x=3
...
3). 3x=6
Divide both sides by 3
x= 2
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com